A box contains 6 defective bulbs and 10 functional bulbs. If 4 bulbs are to be chosen at random. what is the probability that there are 2 defective bulbs and 2 functional bulbs selected?


1
Expert's answer
2022-04-25T12:18:40-0400

There are "16" different bulbs. There are "C_{16}^4=\frac{16!}{4!12!}=\frac{13\cdot14\cdot15\cdot16}{2\cdot3\cdot4}=1820" ways to choose "4" bulbs. There are "C_6^2=15" ways to make a set that contains "2" defective bulbs and "C_{10}^2=45" ways to make a set that contains "2" functional bulbs. We get the following probability: "p=\frac{C_6^2C_{10}^2}{C_{16}^4}=\frac{15\cdot45}{1820}=\frac{135}{364}\approx0.3709".

Thus, the answer is "0.3709" (it is rounded to "4" decimal places).